OpenStudy (anonymous):

If sin θ = 4 5 and tan θ = − 4 3 , find sec θ.

OpenStudy (perl):

sin t = 4/5 tan t = 4/3 find sec t ?

OpenStudy (anonymous):

I need to find sec theta...apparently my copy and paste didnt work so well...it is if sin theta 4/5 and tan theta is -4/3 , find sec theta